What coefficient of friction is needed to keep the
Answer #1

Let the Coefficient of kinetic friction be uk .

Tension in the string, T = 3.0 * 9.8 = 29.4 N

For Block A not to slide,
Fr = T
uk*5.0*9.8 = 29.4
uk = 0.6